Suppose that the table below represents the results of a survey of 250 residents in an...

  1. Suppose that the table below represents the results of a survey of 250 residents in an apartment  that were asked whether they prefer chips, cookies, or fruit and whether they were pregnant or not. Let A denote the event that a person selected at random prefers chips and B denote the event that a person selected at random is pregnant.
    1. Calculate P(A)
    2. Calculate P(B)
    3. Calculate P(A and B)
    4. Calculate P(A or B)
    5. Calculate P (B - A)
    6. Calculate P (A|B)
    7. Calculate P(B|Ac)

Chips

Cookies

Fruit

Total

Pregnant

54

23

42

119

Not pregnant

22

11

98

131

Total

76

34

140

250
Answer #1

Let A denote the event that a person selected at random prefers chips and B denote the event that a person selected at random is pregnant.

Hence, n(A) = Number of person prefer chips = 76. n(B) = number of person is pregnant = 119,

n(S) = total number of person. = 250

a) P(A) = n(A) / n(S) = 76/250 = 0.304

b) P(B) = n(B) / n(S) = 119 / 250 = 0.476

c) P( A and B) = n(A and B) / n(S) = 54/250 = 0.216

d) P(A or B) = P(A) + P(B) - P(A and B) = 0.304+0.476-0.216 = 0.564

e) P(B - A) = P(B) - P(A and B) = 0.476 - 0.216 = 0.260

f) P(A |B) = 54/119 = 0.454

g) P(B | Ac ) = P( B and Ac ) / P(Ac ) = P(B-A) / (1-P(A)) = 0.260 / (1-0.304) = 0.855

****If you have any queries or doubts please comment below. if you're satisfied please give a like. Thank you!